### v2

@@ I Introduction
In this paper, we will continue to discuss the topic of detecting $k$-nonseparability and $k$-partite entanglement based on MUMs and the generalized Wigner-Yanase skew information. The structure of this paper is as follows. In Section 2, we give some background knowledge on several quantum informations, $k$-separability, and $k$-partite entanglement in multipartite quantum systems. Section 3 is devoted to derive criteria for detecting $k$-nonseparability and $k$-partite entanglement in multipartite quantum systems by combining MUMs and the generalized Wigner-Yanase skew information. In addition, we provide examples to demonstrate the effectiveness of these criteria as well as an application to distinct quantum networks. Section 4 is a conclusion.
@@ II Preliminaries
Throughout this paper, assume that $H_{i}$ is any complex Hilbert space with $\dim H_{i}=d_{i}<\infty$, $i=1,2,\cdots,N$, and let $H=H_{1}\otimes H_{2}\otimes\cdots\otimes H_{N}$. Denote by $\mathcal{B}(H)$ the algebra of all bounded linear operators and ${\mathcal{S}}(H)=\{\rho:\rho\geq 0,{\rm Tr}(\rho)=1\}$ is the closed convex set of all quantum states in $\mathcal{B}(H)$. $\mathbb{I}\in{\mathcal{B}}(H)$ and $\mathbb{I}_{i}\in{\mathcal{B}}(H_{i})$ stand for the identity operator on $H$ and $H_{i}$, respectively.
@@ II.1 $k$-nonseparability and $k$-partite entanglement
where $\gamma_{1}|\gamma_{2}|\cdots|\gamma_{m}$ is a $m$-partition of $\{1,2,\cdots,N\}$, and the number of particles in each subset $\gamma_{i}$ is not greater than $k$. For any mixed state $\rho$, we say that $\rho$ is $k$-producible if it can be represented as a convex combination of $k$-producible pure states; otherwise, we say that $\rho$ contains $(k+1)$-partite entanglement GOTG .
@@ II.5 Mutually unbiased measurements (MUMs)
Let $P^{(b)}=\{P_{n}^{(b)}:n=1,\cdots,d\}$ be two positive-operator-valued measurements (POVMs), $b=1,2$. $P^{(1)}$ and $P^{(2)}$ are called MUMs if they satisfy the following conditions:
@@
Kalev and Gour in KG proposed a construction method for MUMs in any dimensional quantum system. Take an orthonormal basis $\{F_{k}\}_{k=1}^{d^{2}-1}$ consisting of Hermitian and traceless operators acting on $H$. Arrange these basis elements on a grid of d1d-1 columns and $d+1$ rows, and denote them as $\{F_{n,b}\}$ with $n=1,2,\ldots,d-1$ and $b=1,2,\ldots,d+1$. Define $d(d+1)$ operators $\{F_{n}^{(b)}\}$ by
@@
The measurement operators $\{P_{n}^{(b)}\}$ are given by:
$P_{n}^{(b)}=\frac{1}{d}\mathbb{I}+tF_{n}^{(b)},$ (3)
where the free parameter $t$ is chosen such that $P_{n}^{(b)}\geq 0$; and
$\kappa=\frac{1}{d}+t^{2}(1+\sqrt{d})^{2}(d-1).$ (4)
Then $\{P^{(b)}=\{P_{n}^{(b)}\}\}$ form a complete set of $d+1$ mutually unbiased measurements, where $b$ labels the measurement and $n$ labels the outcome. Moreover, any complete set of MUMs has this form.
It was also proven in LL that $\left\{\frac{\mathbb{I}}{\sqrt{d}},F_{n,b}:n=1,\cdots,d-1;b=1,\cdots,d+1\right\}$ forms a set of local orthogonal observables (LOOs) on $H$; and any complete set of MUMs $\{P_{n}^{(b)}:n=1,2,\cdots,d;b=1,2,\cdots,d+1\}$ must satisfy the following equation:
@@ III $k$-separability and $k$-partite entanglement detection criteria
For convenience, assume $H_{i}=H_{0}$ with $\dim H_{0}=d$, $i=1,2,\cdots,N$. Take the same set of $d+1$ MUMs acting on each subsystem $H_{i}$
$\{P^{(i,b)}=\{P_{i,n}^{(b)}\}:b=1,2,\cdots,d+1;n=1,2,\cdots,d\},$
@@
$\leq$ $\frac{d+1}{d}(\otimes_{i\in\Gamma}\mathbb{I}_{i})+t^{2}(d+\sqrt{d})^{2}(1-\frac{1}{d})(\otimes_{i\in\Gamma}\mathbb{I}_{i})=(1+\kappa)(\otimes_{i\in\Gamma}\mathbb{I}_{i}).$
$\Box$
@@
Since $\rho$ is $k$-separable, then $\rho$ can be written as a convex combination of $k$-separable pure states, that is, $\rho=\sum_{r}p_{r}\rho_{r}$, where $\rho^{(r)}$ has the form
@@
It is obvious that $y_{0}\in X$. Recall that, for any vectors $x=(x_{1},x_{2},\cdots,x_{k})^{t}$ and $y=(y_{1},y_{2},\cdots,y_{k})^{k}\in{\mathbb{R}}^{k}$, $x\prec y$ means that $\sum_{i=1}^{m}x_{i}\leq\sum_{i=1}^{m}y_{i}$ holds for all $m=1,2,\cdots,k-1$ and $\sum_{i=1}^{k}x_{i}=\sum_{i=1}^{k}y_{i}$. Then we have $x\prec y_{0}$ for all $x\in X$; otherwise, if there exists some $m^{\prime}$ such that $\sum_{i=1}^{m^{\prime}}x_{i}>\sum_{i=1}^{m^{\prime}}y_{i}=N-k+m^{\prime},$ then there must be some $x_{0}=0$, a contradiction. Note that the function $x\mapsto\sum_{i=1}^{k}x_{i}^{2}$ is Schur convex. It follows that $\sum_{i=1}^{k}x_{i}^{2}\leq(N-k+1)^{2}+(k-1)$. This implies that $\sum_{i=1}^{k}\sharp(\Gamma_{i})^{2}$ attains the maximum at the partition $\{\Gamma_{1},\cdots,\Gamma_{k}\}$ with $\sharp(\Gamma_{1})=N-k+1,\ \sharp(\Gamma_{i})=1$ for all $i=2,3,\cdots,k$.
@@
$\begin{array}[]{rl}&\sum_{b=1}^{d+1}\sum_{n=1}^{d}I^{s}(\rho,{\bf P}_{N,n}^{(b)})\\ &\leq pk^{2}(\kappa-\frac{1}{d})+N(d\kappa-1)=Nk(\kappa-\frac{1}{d})+N(d\kappa-1);\end{array}$
@@
$\begin{array}[]{rl}&\sum_{b=1}^{d+1}\sum_{n=1}^{d}I^{s}(\rho,{\bf P}_{N,n}^{(b)})\\ &\leq pk^{2}(\kappa-\frac{1}{d})+(N-pk)^{2}(\kappa-\frac{1}{d})+N(d\kappa-1)\\ &=(N^{2}+p^{2}k^{2}+pk^{2}-2pkN)(\kappa-\frac{1}{d})+N(d\kappa-1).\end{array}$
@@
Consider the 6-partite quantum state
@@
where $|\psi\rangle=\frac{1}{\sqrt{6}}(|000001\rangle+|000010\rangle+|000100\rangle+|001000\rangle+|010000\rangle+|100000\rangle)$.
@@
Consider an 11-partite quantum state
$\rho=p|\psi\rangle\langle\psi|+\frac{1-p}{2^{11}}\mathbb{I}$
where $|\psi\rangle=\frac{1}{\sqrt{2}}(|0\rangle^{\otimes 11}+|1\rangle^{\otimes 11})$.
In this example, we take $s=$-$\infty$ for an illustration. According to Theorem 3 and Theorem 4, we can conclude that $\rho(p)$ is $k$-nonseparable if $p_{k}<p\leq 1$ and exhibits $(k+1)$-partite entanglement if $\tilde{p_{k}}<p\leq 1$. Note that, $p_{k}^{\prime}$ is the result obtained using the method in HLS , and $\tilde{p_{k}^{\prime}}$ is obtained in HQGY , as shown in the Tables VII-VIII.
@@
A quantum network is a system that connects quantum devices to share quantum information, enabling tasks like secure communication and distributed computing using quantum properties LMX ; KHJ . Generally, a quantum network consists of $n$ nodes (subsystems, denote by $A_{j}$, $j=1,2,\cdots,n$) and $r$ quantum resources $\rho_{i}$ shared by some of the nodes. Let $L_{i}\subseteq\{1,2,\cdots,n\}$ the subset of index of the nodes that $\{A_{j}:j\in L_{i}\}$ is exact the subset of all nodes that share the $i$-th resource $\rho_{i}$. Denote by $\rho_{i}=\rho_{L_{i}}$, $i=1,2,\cdots,r$. Then the corresponding quantum network state has the form:
$\rho=\rho_{L_{1}}\otimes\rho_{L_{2}}\otimes\cdots\rho_{L_{r}}.$
Accordingly, for every $j=1,2,\ldots,n$, the node $A_{j}$ may share several $\rho_{L_{i}}$ with other nodes, that is $j\in L_{i}$. Let $F_{j}=\{i_{j}:A_{j}\in L_{i}\}\subseteq\{1,2,\ldots,r\}$. Then the subsystem hold by $A_{j}$ is a $\sharp F_{j}$-partite composite system $H_{A_{j}}=\bigotimes_{i_{j}\in F_{j}}H_{A_{j_{i}}}$. Every quantum network state determined by this network has the form:
$\rho=\sum_{\lambda\in\Lambda}p_{\Lambda}(\lambda)(\Omega_{1}^{\lambda}\otimes\cdots\otimes\Omega_{n}^{\lambda})(\rho_{L_{1}}^{\lambda}\otimes\rho_{L_{2}}^{\lambda}\otimes\cdots\rho_{L_{r}}^{\lambda}),$ (6)
where $\{p_{\Lambda}(\lambda)\}$ is a probability distribution, $\Omega_{j}^{\lambda}$ is a quantum channel of the subsystem $A_{j}$ NWRP . So the number $n$ of nodes corresponds to the number of components in the composite system involved in the network, the number $r$ of factors contained in the product state corresponds to the number of resources. The number of nodes involved in the resource with the most shared nodes $k=\sharp(L)=\max\limits_{i}\sharp(L_{i})$ as the depth of the network. It reflects the fact that the quantum network state with respect to this network can be obtained by preparing the states in at most $k$-partite system, and the corresponding network states of the form in Eq.(6) are called network states with depth $k$.
@@
In this example, we also take $s=$-$\infty$ for an illustration. To do this, we observe that, for the network net 1 in Fig.1, parties $A,B$ and $C$ share a resource state $\rho_{1}$, while parties $C,D$ and $E$ share a resource state $\rho_{2}$. This indicates that the depth of the network net 1 is 3 and the corresponding network state must be 3-producible. Similarly, the depth of the network net 2 is 2 and the corresponding network state is 2-producible, while the depth of the network net 3 is 4 and the corresponding network state is 4-producible.
@@
Then, by Theorem 4, it is easily seen that $\rho_{c}$ is not 3-producible and thus is the network state corresponding to the network of depth greater than 3. Consequently, $\rho_{c}$ is the state corresponding to net 3. Similarly, one concludes that $\rho_{b}$ is the network state corresponding to the network net 1 and $\rho_{a}$ is corresponding to net 2. $\Box$

### v3

@@ I Introduction
In this paper, we will continue to discuss the topic of detecting $k$-nonseparability and $k$-partite entanglement based on MUMs and the generalized Wigner-Yanase skew information. The structure of this paper is as follows. In Section 2, we give some background knowledge on quantum informations, $k$-separability, and $k$-partite entanglement in multipartite quantum systems. Section 3 is devoted to derive criteria for detecting $k$-nonseparability and $k$-partite entanglement in multipartite quantum systems by combining MUMs and the generalized Wigner-Yanase skew information. In addition, we provide examples to demonstrate the effectiveness of these criteria as well as an application to distinct quantum networks. Section 4 is a conclusion.
@@ II Preliminaries
Throughout this paper, assume that $H_{i}$ is any complex Hilbert space with $\dim H_{i}=d_{i}<\infty$, $i=1,2,\cdots,N$, and let $H=H_{1}\otimes H_{2}\otimes\cdots\otimes H_{N}$. Denote by $\mathcal{B}(H)$ the algebra of all bounded linear operators and ${\mathcal{S}}(H)=\{\rho:\rho\geq 0,{\rm Tr}(\rho)=1\}$ is the closed convex set of all quantum states in $\mathcal{B}(H)$. $\mathbb{I}\in{\mathcal{B}}(H)$ and $\mathbb{I}_{i}\in{\mathcal{B}}(H_{i})$ stand for the identity operators on $H$ and $H_{i}$, respectively.
@@ II.1 $k$-nonseparability and $k$-partite entanglement
where $\gamma_{1}|\gamma_{2}|\cdots|\gamma_{m}$ is a $m$-partition of $\{1,2,\cdots,N\}$, and the number of particles in each subset $\gamma_{i}$ is not greater than $k$. For any mixed state $\rho$, we say that $\rho$ is $k$-producible if it can be represented as a convex combination of $k$-producible pure states; otherwise, we say that $\rho$ contains $(k+1)$-partite entanglement, or is $(k+1)$-partite entangled GOTG .
@@ II.5 Mutually unbiased measurements (MUMs)
Let $P^{(b)}=\{P_{n}^{(b)}:n=1,\cdots,d\}\ (b=1,2)$ be two positive-operator-valued measurements (POVMs) on a $d$-dimensional Hilbert space $H_{0}$. $P^{(1)}$ and $P^{(2)}$ are called MUMs if they satisfy the following conditions:
@@
Kalev and Gour in KG proposed a construction method for MUMs in any dimensional quantum system. Take an orthonormal basis $\{F_{k}\}_{k=1}^{d^{2}-1}$ for $B_{0}(H_{0})$, consisting of Hermitian and traceless operators acting on $H_{0}$ and with inner product $(A,B)={\rm Tr}(B^{\dagger}A)$. Arrange these basis elements on a grid of d1d-1 columns and $d+1$ rows: the first row consists of $F_{1},F_{2},\ldots,F_{d-1}$, the second row consists of $F_{d},F_{d+1},\ldots,F_{2(d-1)}$, and so on. Denote the element in the $n$-th column of the $b$-th row by $F_{n,b}$, with $n=1,2,\ldots,d-1$ and $b=1,2,\ldots,d+1$. Define $d(d+1)$ operators $F_{n}^{(b)}$ by
@@
It was proved in KG that there exists a parameter $t$, called the free parameter, such that the operators $P_{n}^{(b)}$ given by
$P_{n}^{(b)}=\frac{1}{d}\mathbb{I}+tF_{n}^{(b)}$ (3)
are positive for all $n,b$. Put
$\kappa=\frac{1}{d}+t^{2}(1+\sqrt{d})^{2}(d-1),$ (4)
then $P^{(b)}=\{P_{n}^{(b)}\}_{n=1}^{d}$ ($b=1,2,\cdots,d+1$) form a complete set of $d+1$ MUMs, where $b$ labels the measurement and $n$ labels the outcome. Moreover, any complete set of MUMs has this form and satisfies the following equation:
@@
It was also shown in $L$ that
$\left\{\frac{\mathbb{I}}{\sqrt{d}},F_{n,b}\colon n=1,\dots,d-1;\,b=1,\dots,d+1\right\}$
forms a set of local orthogonal observables (LOOs) on $H_{0}$.
@@ III $k$-separability and $k$-partite entanglement detection criteria
For convenience, assume $H_{i}=H_{0}$ with $\dim H_{0}=d$, $i=1,2,\dots,N$. Take the same $d+1$ MUMs on each subsystem $H_{i}$:
$P^{(i,b)}=\{P_{i,n}^{(b)}\}_{n=1}^{d}\ \ (b=1,2,\cdots,d+1),$
@@
$\leq$ $\frac{d+1}{d}(\otimes_{i\in\Gamma}\mathbb{I}_{i})+t^{2}(d+\sqrt{d})^{2}(1-\frac{1}{d})(\otimes_{i\in\Gamma}\mathbb{I}_{i})$
$=$ $(1+\kappa)(\otimes_{i\in\Gamma}\mathbb{I}_{i}).$
Here, the first equality uses Eq.(3), the third equality uses Eq.(2), the fourth equality uses Eq.(1) and the last equality uses Eq.(4). $\Box$
@@
Since $\rho$ is $k$-separable, $\rho$ can be written as a convex combination of $k$-separable pure states, that is, $\rho=\sum_{r}p_{r}\rho_{r}$, where $\rho^{(r)}$ has the form
@@
It is obvious that $y_{0}\in X$. Recall that, for any nonnegative vectors $x=(x_{1},x_{2},\cdots,x_{k})^{t}$ and $y=(y_{1},y_{2},\cdots,y_{k})^{t}\in{\mathbb{R}}^{k}$ with coordinate descending order arrangement, $x\prec y$ means that $\sum_{i=1}^{m}x_{i}\leq\sum_{i=1}^{m}y_{i}$ holds for all $m=1,2,\cdots,k-1$ and $\sum_{i=1}^{k}x_{i}=\sum_{i=1}^{k}y_{i}$. Then we have $x\prec y_{0}$ for all $x\in X$; otherwise, if there exists some $m^{\prime}$ such that $\sum_{i=1}^{m^{\prime}}x_{i}>\sum_{i=1}^{m^{\prime}}y_{i}=N-k+m^{\prime},$ then there must be some $x_{0}=0$, a contradiction. Note that the function $x\mapsto\sum_{i=1}^{k}x_{i}^{2}$ is Schur convex. It follows that $\sum_{i=1}^{k}x_{i}^{2}\leq(N-k+1)^{2}+(k-1)$. This implies that $\sum_{i=1}^{k}\sharp(\Gamma_{i})^{2}$ attains the maximum at the partition $\{\Gamma_{1},\cdots,\Gamma_{k}\}$ with $\sharp(\Gamma_{1})=N-k+1,\ \sharp(\Gamma_{i})=1$ for all $i=2,3,\cdots,k$.
@@
$\begin{array}[]{rl}&\sum_{b=1}^{d+1}\sum_{n=1}^{d}I^{s}(\rho,{\bf P}_{N,n}^{(b)})\\ \leq&pk^{2}(\kappa-\frac{1}{d})+N(d\kappa-1)=Nk(\kappa-\frac{1}{d})+N(d\kappa-1);\end{array}$
@@
$\begin{array}[]{rl}&\sum_{b=1}^{d+1}\sum_{n=1}^{d}I^{s}(\rho,{\bf P}_{N,n}^{(b)})\\ \leq&pk^{2}(\kappa-\frac{1}{d})+(N-pk)^{2}(\kappa-\frac{1}{d})+N(d\kappa-1)\\ =&(N^{2}+p^{2}k^{2}+pk^{2}-2pkN)(\kappa-\frac{1}{d})+N(d\kappa-1).\end{array}$
@@
We remark here that, although the supremum over $s\in[-\infty,0]$ in Theorems 3-4 is theoretically difficult to evaluate due to the nonlinear dependence of $\sum_{b=1}^{d+1}\sum_{n=1}^{d}I^{s}(\rho,{\bf P}_{N,n}^{(b)})$ on $s$, in practice it can be well approximated by considering extreme values such as $s=$-$\infty$ or by sampling $s$ at a finite set of points. This allows effective detection of $k$-nonseparability while retaining the practical advantages of the criterion.
@@
Consider the 6-partite quantum state on the Hilbert space $H=(\mathbb{C}^{2})^{\otimes 6}$:
@@
where $|\psi\rangle=\frac{1}{\sqrt{6}}(|000001\rangle+|000010\rangle+|000100\rangle+|001000\rangle+|010000\rangle+|100000\rangle)$ and $0\leq p\leq 1$.
@@
Consider an 11-partite quantum state on the Hilbert space $H=(\mathbb{C}^{2})^{\otimes 11}$:
$\rho=p|\psi\rangle\langle\psi|+\frac{1-p}{2^{11}}\mathbb{I},$
where $|\psi\rangle=\frac{1}{\sqrt{2}}(|0\rangle^{\otimes 11}+|1\rangle^{\otimes 11})$ and $0\leq p\leq 1$.
In this example, we take $s=$-$\infty$ for an illustration. According to Theorem 3 and Theorem 4, we can conclude that $\rho(p)$ is $k$-nonseparable if $p_{k}<p\leq 1$ and exhibits $(k+1)$-partite entanglement if $\tilde{p_{k}}<p\leq 1$. Note that, $p_{k}^{\prime}$ is the result obtained using the method in HLS , and $\tilde{p_{k}^{\prime}}$ is obtained in HQGY2 , as shown in the Tables VII-VIII.
@@
A quantum network is a system that connects quantum devices to share quantum information, enabling tasks like secure communication and distributed computing using quantum properties LMX ; KHJ . Generally, a quantum network consists of $n$ nodes (subsystems, denote by $A_{j}$, $j=1,2,\cdots,n$) and $r$ quantum resources $\rho_{i}$ shared by some of the nodes. Let $L_{i}\subseteq\{1,2,\cdots,n\}$ be the subset of index of the nodes that $\{A_{j}:j\in L_{i}\}$ is exact the subset of all nodes that share the $i$-th resource $\rho_{i}$. Denote by $\rho_{i}=\rho_{L_{i}}$, $i=1,2,\cdots,r$. Then the corresponding quantum network state has the form:
$\rho=\rho_{L_{1}}\otimes\rho_{L_{2}}\otimes\cdots\otimes\rho_{L_{r}}.$
Accordingly, for every $j=1,2,\ldots,n$, the node $A_{j}$ may share several $\rho_{L_{i}}$ with other nodes, that is $j\in L_{i}$. Let $F_{j}=\{i_{j}:A_{j}\in L_{i}\}\subseteq\{1,2,\ldots,r\}$. Then the subsystem hold by $A_{j}$ is a $\sharp F_{j}$-partite composite system $H_{A_{j}}=\bigotimes_{i_{j}\in F_{j}}H_{A_{i_{j}}}$. Every quantum network state determined by this network has the form:
$\rho=\sum_{\lambda\in\Lambda}p_{\Lambda}(\lambda)(\Omega_{1}^{\lambda}\otimes\cdots\otimes\Omega_{n}^{\lambda})(\rho_{L_{1}}^{\lambda}\otimes\rho_{L_{2}}^{\lambda}\otimes\cdots\otimes\rho_{L_{r}}^{\lambda}),$ (6)
where $\{p_{\Lambda}(\lambda)\}$ is a probability distribution, $\Omega_{j}^{\lambda}$ is a quantum channel of the subsystem $A_{j}$ NWRP . So the number $n$ of nodes corresponds to the number of components in the composite system involved in the network, the number $r$ of factors contained in the product state corresponds to the number of sources. Denote by $\sharp(L_{i})$ the number of nodes involved in the source $\rho_{i}$. The most shared nodes $k=\max\limits_{i}\sharp(L_{i})$ is called the depth of the network, which reflects the fact that the quantum network state with respect to this network can be obtained by preparing the states in at most $k$-partite system, and the corresponding network states of the form in Eq.(6) are called network states with depth $k$.
@@
In this example, we also take $s=$-$\infty$ for an illustration. To do this, we observe that, for the network net 1 in Fig.1, parties $A,B$ and $C$ share a source state $\rho_{1}$, while parties $C,D$ and $E$ share a source state $\rho_{2}$. This indicates that the depth of the network net 1 is 3 and the corresponding network state must be 3-producible. Similarly, the depth of the network net 2 is 2 and the corresponding network state is 2-producible, while the depth of the network net 3 is 4 and the corresponding network state is 4-producible.
@@
Then, by Theorem 4, it is easily seen that $\rho_{c}$ is not 3-producible and thus is the network state corresponding to the network of depth greater than 3. Consequently, $\rho_{c}$ is the state corresponding to net 3. Similarly, one concludes that $\rho_{b}$ is the network state corresponding to the network net 1 and $\rho_{a}$ is corresponding to net 2. This completes the given task. $\Box$